OpenStudy (anonymous):

Friends plse tell how to solve it What must be subtracted from pq+q^2 to get -6p^2+15p^2-pq

OpenStudy (anonymous):

actually i forgot Number- pq=-6p^2+15p^2-pq or pq-number=-6p^2+15p^2-pq

OpenStudy (anonymous):

here number is to be find out

OpenStudy (anonymous):

its the second option! pq+q^2 - number = -6p^2+15p^2-pq and then number = pq+q^2 - (-6p^2+15p^2-pq)

OpenStudy (anonymous):

@Soumyadeep thankk u so much